OpenStudy (anonymous):

Enter the value of the question mark as a fraction: ^4 sqrt(5^) = 5^?

OpenStudy (anonymous):

\[\sqrt[4]{5^5} = 5^{?}\]

OpenStudy (anonymous):

^ Proper format!

OpenStudy (misty1212):

\[\huge \sqrt[\color{red}m]x^{\color{green}n}=x^{\frac{\color{green}n}{\color{red}m}}\]